# One-Shot Real-to-Sim via End-to-End Differentiable Simulation and Rendering

### v3

@@ V-B Physical Experiments
Shown in Fig. 3, we conduct physical experiments with a UR5e robot arm equipped with an embedded end-effector F/T sensor and a pusher, and a static overhead RealSense D435 RGB-D camera. We use one training trajectory and three testing pushing trajectories that all share the same starting location but are spaced about 15 apart as the simulated experiments for 3 YCB objects: power drill (Drill), sugar box (Sugar), and mustard bottle (Mustard). The trajectories push the object up to 9 cm and 30. We use a total of $T=48$ time steps with $\delta t=0.03$ s. We use the same optimization settings as the simulated experiments.
On average across all testing trajectories and testing objects, our method achieves a mass identification error of 0.257 kg and 5.66 mm of unilateral chamfer distance between the observed object point cloud and the predicted object at the last frame of the trajectory. Note that we only reported the mass identification error because the ground truth is very easy to measure while measuring the surface friction requires a specialized setup. Since we do not have access to the ground-truth object poses and only have access to the raw observations, we report the unilateral chamfer distances between the observed object point cloud and the predicted object geometry. The train and test prediction results for three sample trajectories are visualized in Fig. 7. Overall, the errors are comparable to those from the simulated experiments. In addition, we show the initial and optimized shape of the Sugar object in Fig. 8 on the right, where the initial shape is too thin and does not contact the terrain below. Our algorithm modifies the occluded geometry to satisfy the physics.

### v4

@@ V-B Physical Experiments
Shown in Fig. 3, we conduct physical experiments with a UR5e robot arm equipped with an ATI Gamma F/T sensor and a pusher with a semispherical end, and a static overhead RealSense D435 RGB-D camera. We use similar pushing trajectories to those for the simulation experiments, but only use 6 trajectories with two different starting locations from one side of the object. We then use one trajectory for training and the rest for evaluation with 3 YCB objects: power drill (Drill), sugar box (Sugar), and mustard bottle (Mustard). We use a total of $T=48$ time steps with $\delta t=0.03$ s. We use the same optimization settings as the simulated experiments.
On average across all testing trajectories and testing objects, our method achieves a mass identification error of 0.186 kg and 6.10 mm of unilateral chamfer distance between the observed object point cloud and the predicted object at the last frame of the trajectory. Note that we only reported the mass identification error because the ground truth is very easy to measure while measuring the surface friction requires a specialized setup. Since we do not have access to the ground-truth object poses and only have access to the raw observations, we report the unilateral Chamfer distances between the observed object point cloud and the predicted object geometry. The train and test prediction results for three sample trajectories are visualized in Fig. 7. Overall, the errors are comparable to those from the simulated experiments. In addition, we show the initial and optimized shape of the Sugar object in Fig. 8 on the right, where the initial shape is too thin and does not contact the terrain below. Our algorithm modifies the occluded geometry to satisfy the physics.